. Determine the force (per unit area) which two like-charged infinitely long planes having the same charge surface density of 3 x 10^-8 C/cm2 will repel each other with


1
Expert's answer
2022-01-30T13:52:41-0500

Answer

charge surface density of

"\sigma=3 *10^{-4 }C/m^2"

the force (per unit area) which two like-charged infinitely long planes

Is given by

"\frac{F}{A}=\frac{\sigma^2}{2\epsilon_0}"

Putting all values


"\frac{F}{A}=\frac{ (3 *10^{-4 } ) ^2}{2\times8.85*10^{-12}}"

"=5084.7N\/m^2"